# Optimized Speculative Sampling for GPU Hardware Accelerators

### v1

@@ 4.2 Main results
Table 1 summarizes accuracy metrics and profiling results for ASR and text summarization tasks. The table details the datasets, target and draft models used, performance metrics, and the relative reduction in overall GPU profiling time achieved by our optimized approaches (exact and sigmoid approximation) compared to the baseline.
In the ASR task, our exact optimization method maintains the same WER compared to the baseline and achieves reduction in profiling time ranging from 8.7% to 12.5%. The sigmoid approximation approach results in slightly increased WERs, but yields more significant profiling time improvements, ranging from 71.9% to 84.0%.
@@
Additionally, we provide relative wall-clock time improvements for the overall text generation process in Table 3 of Appendix § A.5, showing that the results obtained via exact profiling translate into overall improvements in wall-clock time for both the exact and the sigmoid approximation method.
@@ 4.3 Analysis and discussion
We assess the memory usage of our optimized methods relative to the baseline implementation. Fig. 4 shows the peak memory usage (HBM) on randomly sampled instances of the Xsum test set with various initial values of $\gamma$ and different language models. The graph indicates that our optimized approaches do not introduce additional memory overhead compared to the baseline. For all three model combinations (Llama2, Qwen, and Gemma), the memory usage of the optimized methods fluctuates slightly (within a range of approximately 200MB) around the memory usage of the baseline across all draft tokens. As illustrated in Fig. 5, the memory usage results for the Whisper models display a pattern consistent with the text summarization experiments, showing fluctuation within a range of under 10 MB.
@@
To study the effect of logit scaling in the sigmoid approximation method, we compare profiling time and performance metrics under varying values of $\alpha$ and $\beta$ (cf. Eq. 5). Table 2 provides a comparison of different scaling factors for subsets (10%) of CV16 and Xsum using Whisper Small and Llama2 7B, respectively. For each task, the table shows the values of $\alpha$ and $\beta$ applied, the resulting WER or ROUGE-1 score, and the relative improvement in profiling time over the non-optimized baseline implementation. Note that scaling is necessary due to the numerical instability induced by the exponentiation in sigmoid.
The Llama2 model combination exhibits relative stability across different scaling factors, showing minor fluctuations in ROUGE-1 scores and profiling time improvements, whereas the Whisper model combination is more sensitive, with scaling factors of ±$\pm 10^{5}$ leading to substantial deterioration of both WER and profiling time. This can be attributed to the logits of Whisper models being generated in half precision, whereas the logits generated by the Llama models are available in full precision. However, we also find that scaling factors of ±$\pm 10^{3}$ and ±$\pm 10^{4}$ generally yield comparable results in both accuracy and profiling time improvement across the model combinations investigated in this work. The results of the same analysis for the other draft and target model combinations are in Table 5 from Appendix § A.6.
@@
Draft/Target Scale ($\alpha,\beta$) WER ()($\downarrow$)( ↓ ) $\Delta$Prof. Time
Whisper Small.EN Distil Small.EN Baseline 0.24
@@
Draft/Target Scale ($\alpha,\beta$) ROUGE-1 ()($\uparrow$)( ↑ ) $\Delta$Prof. Time
Llama2 7B/ Sheared 1.3B Baseline 0.19
@@ 5 Conclusions
We introduced two optimization methods to accelerate speculative sampling for autoregressive models on hardware accelerators. By computing significant portions of intermediate matrices across multiple GPU threads within thread blocks and by minimizing slow read and write operations across the memory hierarchy, our exact optimization method led to improved sampling speed without compromising accuracy. Additionally, we employed an approximation technique using element-wise sigmoid instead of softmax, to enable parallel computation of probabilities. This approximation further accelerated the decoding process but resulted in a small degradation of sampling quality.
6 Limitations
In this work, we study the inference efficiency of speech and language models in the context of speculative decoding using GPU hardware accelerators. Our investigation includes two optimized speculative sampling algorithms, tested on these models to enhance inference speed. While GPUs are the most common general-purpose hardware accelerators, there exist purpose-built architectures such as Cerebras’s Wafer Scale Engines, Google’s TPUs, and GraphCore’s IPUs, where the differences in system design may negate or significantly reduce our latency gains. A key component of our approach relies on minimizing read and write operations between HBM and SRAM. If this bottleneck is alleviated by the underlying hardware itself, e.g., through a much larger SRAM, our approach is unlikely to yield significant improvements. Additionally, our study evaluates the effectiveness of the optimized algorithm based on decoding time, and our claims may not translate to other metrics, such as energy usage or heat generation, although they play an important role in real-world production settings.
@@ A.2 Draft model details
- - Target Draft Exact Sigmoid
TED-LIUM release3 Small.EN Distil Small.EN 8.7% 58.7%
TED-LIUM release3 Large V2 Distil Large V2 6.6% 66.3%
LibriSpeech clean Small.EN Distil Small.EN 5.8% 49.5%
LibriSpeech clean Large V2 Distil Large V2 7.4% 65.3%
LibriSpeech other Small.EN Distil Small.EN 7.1% 54.5%
LibriSpeech other Large V2 Distil Large V2 7.0% 66.2%
CV16 en Small.EN Distil Small.EN 8.1% 51.8%
CV16 en Large V2 Distil Large V2 5.4% 61.3%
CNN/DailyMail - Gemma 7B Gemma 2B 3.0% 24.2%
CNN/DailyMail - Qwen 7B Qwen 0.5B 1.6% 39.8%
CNN/DailyMail - Llama2 7B Sheared Llama 1.3B 4.2% 28.2%
CNN/DailyMail - Llama2 13B Sheared Llama 1.3B 2.6% 23.5%
XSum - Gemma 7B Gemma 2B 1.2% 18.5%
XSum - Qwen 7B Qwen 0.5B 4.3% 59.1%
XSum - Llama2 7B Sheared Llama 1.3B 6.5% 53.1%
XSum - Llama2 13B Sheared Llama 1.3B 10.9% 23.6%
@@ A.5 Average Times per Decoding Step
Dataset Subset Model Avg.$\pm$±Std. (ms) $\Delta$Profiling Time
- - Target Draft Baseline Exact Sigmoid Exact Sigmoid
@@
TED-LIUM release3 Large V2 Distil Large V2 4.37$\pm$0.58 3.88$\pm$0.46 3.62$\pm$1.09 11.2% 17.2%
@@
LibriSpeech clean Large V2 Distil Large V2 4.35$\pm$0.60 3.88$\pm$0.52 3.55$\pm$1.08 10.8% 18.4%
LibriSpeech other Small.EN Distil Small.EN 4.14$\pm$0.84 3.68$\pm$0.68 3.20$\pm$1.07 11.1% 22.7%
LibriSpeech other Large V2 Distil Large V2 4.39$\pm$0.61 3.90$\pm$0.49 3.55$\pm$1.07 11.2% 19.1%
@@
CV16 en Large V2 Distil Large V2 4.37$\pm$0.61 3.92$\pm$0.51 3.62$\pm$1.07 10.4% 17.2%
CNN/DailyMail - Gemma 7B Gemma 2B 6.54$\pm$0.47 5.84$\pm$0.39 4.37$\pm$0.54 10.6% 33.1%
CNN/DailyMail - Qwen 7B Qwen 0.5B 12.01$\pm$1.08 11.20$\pm$1.03 3.34$\pm$0.45 6.8% 72.2%
CNN/DailyMail - Llama2 7B Sheared Llama 1.3B 11.33$\pm$0.94 10.69$\pm$0.83 3.65$\pm$0.62 5.7% 67.8%
CNN/DailyMail - Llama2 13B Sheared Llama 1.3B 3.99$\pm$0.52 3.59$\pm$1.70 3.26$\pm$0.26 10.1% 18.3%
XSum - Gemma 7B Gemma 2B 6.39$\pm$0.50 5.76$\pm$0.42 4.61$\pm$0.55 9.9% 27.9%
XSum - Qwen 7B Qwen 0.5B 11.55$\pm$1.39 10.65$\pm$1.51 3.20$\pm$0.42 7.8% 72.3%
XSum - Llama2 7B Sheared Llama 1.3B 4.66$\pm$0.42 4.14$\pm$2.88 3.64$\pm$0.56 11.1% 21.8%
XSum - Llama2 13B Sheared Llama 1.3B 4.67$\pm$0.41 4.17$\pm$1.70 3.70$\pm$0.44 10.5% 20.7%
Table 4: Average time and standard deviation spent within the speculative sampling algorithm per decoding step. The column “$\Delta$Profiling Time” measures the relative reduction in average time per decoding step (“Baseline” vs. “Exact” and“Sigmoid”). Scaling constants for sigmoid approximation: $\alpha=-10^{3}$ and $\beta=10^{3}$ for ASR, $\alpha=-10^{4}$ and $\beta=10^{4}$ for summarization.
@@ A.6 Effect of Scaling Logits
Draft/Target Scale ($\alpha,\beta$) WER ()($\downarrow$)( ↓ ) $\Delta$Prof. Time
@@
Draft/Target Scale ($\alpha,\beta$) ROUGE-1 ()($\uparrow$)( ↑ ) $\Delta$Prof. Time

### v2

@@ 4.2 Main results
Table 1 summarizes accuracy metrics and profiling results for the ASR and text summarization tasks. The table details the datasets, target and draft models used, performance metrics, and the relative reduction in overall GPU profiling time achieved by our optimized approaches (exact and sigmoid approximation) compared to the baseline.
In the ASR task, our exact optimization method maintains the same WER compared to the baseline and achieves reduction in profiling time ranging from 8.7% to 12.5%. The sigmoid approximation approach results in moderately increased WERs, but yields more significant profiling time improvements, ranging from 71.9% to 84.0%.
@@
Additionally, we provide relative wall-clock time improvements for the overall text generation process in Table 5 of Section A.5, showing that the results obtained via profiling translate into improvements in wall-clock time for both the exact and the sigmoid approximation method.
@@ 4.3 Analysis and discussion
We assess the memory usage of our optimized methods relative to the baseline implementation. Fig. 4 shows the peak memory usage (HBM) on randomly sampled instances (10%) of the Xsum test set with various initial values of $\gamma$ and different language models. The graph indicates that our optimized approaches do not introduce additional memory overhead compared to the baseline. For all three model combinations (Llama2, Qwen, and Gemma), the memory usage of the optimized methods fluctuates slightly (within a range of approximately 200MB) around the memory usage of the baseline across all draft tokens. As illustrated in Fig. 5, the memory usage results for the Whisper models display a pattern consistent with the text summarization experiments, showing fluctuation within a range of under 10 MB.
@@
To study the effect of logit scaling in the sigmoid approximation method, we compare profiling time and performance metrics under varying values of $\alpha$ and $\beta$ (cf. Eq. 5). Table 2 provides a comparison of different scaling factors for subsets (10%) of CV16 and Xsum using Whisper Small.EN and Llama2 7B, respectively. For each task, the table shows the values of $\alpha$ and $\beta$ applied, the resulting WER or ROUGE-1 score, and the relative improvement in profiling time over the non-optimized baseline implementation. Note that scaling is necessary due to the numerical instability induced by the exponentiation in the sigmoid function.
The Llama2 model combination exhibits relative stability across different scaling factors, showing minor fluctuations in ROUGE-1 scores and profiling time improvements, whereas the Whisper model combination is more sensitive, with scaling factors of ±$\pm 10^{5}$ leading to substantial deterioration of both WER and profiling time. This can be attributed to the logits of Whisper models being generated in half precision, whereas the logits generated by the Llama models are available in full precision. However, we also find that scaling factors of ±$\pm 10^{3}$ and ±$\pm 10^{4}$ generally yield comparable results in both accuracy and profiling time improvement across the model combinations investigated in this work. The results of the same analysis for the other draft and target model combinations are provided in Table 7 from  Section A.6.
@@
Target/Draft Scale ($\alpha,\beta$) WER ()($\downarrow$)( ↓ ) $\Delta$Prof. Time
Whisper Small.EN/ Distil-Whisper Small.EN Baseline 0.24
@@
Target/Draft Scale ($\alpha,\beta$) ROUGE-1 ()($\uparrow$)( ↑ ) $\Delta$Prof. Time
Llama2 7B/ Sheared Llama 1.3B Baseline 0.19
@@
Data transfer between HBM and SRAM.
To assess efficiency in terms of data movement between HBM and on-chip SRAM, we compare the realized bandwidths of each implementation. The bandwidth is calculated by dividing the total bytes transferred by the execution time. The number of bytes transferred is derived from the total number of sectors moved between HBM and SRAM, with each sector consisting of 32 bytes. Execution time refers to the duration during which the GPU is actively running a kernel, i.e., when at least one GPU unit is engaged in computation rather than idling, waiting, or stalled. A lower realized bandwidth indicates a reduced communication overhead between HBM and SRAM.
The results in Table 3 show lower memory bandwidth usage for the Qwen and Gemma models with the exact optimization approach compared to the baseline implementation. However, in the case of the Llama2 and Whisper models, despite overall lower realized bandwidths relative to the Qwen and Gemma models, higher bandwidths are observed with the exact optimization compared to their corresponding baseline.
The sigmoid approximation has consistently higher realized bandwidths across all model combinations compared to the baseline. Although the sigmoid optimization approach reduces the overall amount of data transferred, i.e., the total number of bytes moved between HBM and SRAM, due to the element-wise approximation of the softmax function within the sampling kernel, the significantly faster execution times result in higher overall realized bandwidths. However, even the highest realized bandwidths are far below the theoretical HBM bandwidth limit of $\sim$2 TB/s NVIDIA Corporation (2020), indicating that memory transfer is not the limiting factor for performance.
Model Realized Bandwidth
Draft Target Baseline Exact Sigmoid
Whisper Small.EN Distil-Whisper Small.EN 13.98 GB/s 14.16 GB/s 16.33 GB/s
Whisper Large V2 Distil-Whisper Large V2 9.32 GB/s 11.18 GB/s 16.06 GB/s
Qwen 7B Qwen 0.5B 44.99 GB/s 31.65 GB/s 52.14 GB/s
Gemma 7B Gemma 2B 53.69 GB/s 38.51 GB/s 62.99 GB/s
Llama2 7B Sheared Llama 1.3B 24.56 GB/s 27.70 GB/s 30.52 GB/s
Llama2 13B Sheared Llama 1.3B 20.18 GB/s 24.08 GB/s 31.39 GB/s
Table 3: Comparison of realized bandwidths across models and optimization techniques using 100 examples of the XSum test set for Qwen, Gemma, and Llama2 models and 100 examples of the CV16 test set for Whisper models.
Results on RTX 2080 TI GPU.
Table 4 shows an accuracy and profiling time comparison using RTX 2080 TI GPUs. We observe that our optimization methods achieve performance similar to A100 GPUs with relative improvements in profiling time ranging between $\sim$5and $\sim$13with the exact method and between $\sim$62and $\sim$82with sigmoid approximation. The summarization experiment with Qwen uses a 1.8B/0.5B parameter model combination instead of the 7B/0.5B model combination from the main experiments in Table 1, due to the limited HBM size of 11GB available on the RTX 2080 TI series.
Dataset Subset Target/Draft WER ()($\downarrow$)( ↓ ) $\Delta$ Prof. Time
Basel. Exact Sigm. Exact Sigmoid
LibriSpeech clean Whisper Small.EN/ Distil Small.EN 0.08 0.08 0.09 7.6% 66.8%
LibriSpeech other 0.14 0.14 0.15 12.3% 65.9%
TED-LIUM release3 0.22 0.22 0.24 12.6% 62.8%
CV16 en 0.22 0.22 0.27 12.3% 62.0%
LibriSpeech clean Whisper Large V2/ Distil Large V2 0.07 0.07 0.15 11.2% 82.5%
LibriSpeech other 0.12 0.12 0.19 5.7% 79.2%
TED-LIUM release3 0.20 0.20 0.23 9.4% 79.8%
CV16 en 0.25 0.25 0.31 7.8% 75.6%
Dataset Subset Target/Draft ROUGE-1 ()($\uparrow$)( ↑ ) $\Delta$ Prof. Time
Base. Exact Sigm. Exact Sigmoid
Xsum Qwen 1.8B Qwen 0.5 0.15 0.15 0.11 5.0% 76.8%
Table 4: Accuracy and profiling time comparison across various datasets and models using a RTX 2080 TI GPUs.
@@ 5 Conclusions
We introduced two optimization methods to accelerate speculative sampling for autoregressive models on hardware accelerators. By computing significant portions of intermediate matrices across multiple GPU threads within thread blocks, our exact optimization method led to improved sampling speed without compromising accuracy. Additionally, we employed an approximation technique using element-wise sigmoid instead of softmax, to enable parallel computation of probabilities. This approximation further accelerated the decoding process but resulted in a small degradation of sampling quality.
Limitations
In this work, we study the inference efficiency of speech and language models in the context of speculative decoding using GPU hardware accelerators. Our investigation includes two optimized speculative sampling algorithms, tested on these models to enhance inference speed. While GPUs are the most common general-purpose hardware accelerators, there exist purpose-built architectures such as Cerebras’s Wafer Scale Engines, Google’s TPUs, and GraphCore’s IPUs, where the differences in system design may negate or significantly reduce the latency gains. Our experiments were conducted exclusively on A100 and RTX 2080 TI GPUs on a single compute node. Therefore, the generalizability of the results to other hardware configurations remains uncertain. The performance outcomes may be influenced by other hardware and network configurations, such as multi-GPU and multi-node setups, as well as the availability fast interconnects (e.g. Infiniband), and other network conditions. Additionally, our study evaluates the effectiveness of the optimized algorithm based on decoding time, and our claims may not translate to other metrics, such as energy usage or heat generation, although they play an important role in real-world production settings.
Acknowledgments
We gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b196ac14. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. This work was supported by the Bavarian State Ministry of Science and the Arts under grant H.2-F1116.NÜ/61/2.
@@ A.2 Draft model details
Target Draft Exact Sigmoid
TED-LIUM release3 Whisper Small.EN Distil-Whisper Small.EN 8.7% 58.7%
release3 Whisper Large V2 Distil-Whisper Large V2 6.6% 66.3%
LibriSpeech clean Whisper Small.EN Distil-Whisper Small.EN 5.8% 49.5%
clean Whisper Large V2 Distil-Whisper Large V2 7.4% 65.3%
other Whisper Small.EN Distil-Whisper Small.EN 7.1% 54.5%
other Whisper Large V2 Distil-Whisper Large V2 7.0% 66.2%
CV16 en Whisper Small.EN Distil-Whisper Small.EN 8.1% 51.8%
en Whisper Large V2 Distil-Whisper Large V2 5.4% 61.3%
CNN/DailyMail Gemma 7B Gemma 2B 3.0% 24.2%
Qwen 7B Qwen 0.5B 1.6% 39.8%
Llama2 7B Sheared Llama 1.3B 4.2% 18.2%
Llama2 13B Sheared Llama 1.3B 2.6% 23.5%
Xsum Gemma 7B Gemma 2B 1.2% 18.5%
Qwen 7B Qwen 0.5B 4.3% 59.1%
Llama2 7B Sheared Llama 1.3B 6.5% 53.1%
Llama2 13B Sheared Llama 1.3B 10.9% 23.6%
@@ A.5 Average times per decoding step
Dataset Subset Model Avg.$\pm$±Std. (ms) $\Delta$Prof. Time
Target Draft Baseline Exact Sigmoid Exact Sigmoid
@@
release3 Large V2 Distil Large V2 4.37$\pm$0.58 3.88$\pm$0.46 3.62$\pm$1.09 11.2% 17.2%
@@
clean Large V2 Distil Large V2 4.35$\pm$0.60 3.88$\pm$0.52 3.55$\pm$1.08 10.8% 18.4%
other Small.EN Distil Small.EN 4.14$\pm$0.84 3.68$\pm$0.68 3.20$\pm$1.07 11.1% 22.7%
other Large V2 Distil Large V2 4.39$\pm$0.61 3.90$\pm$0.49 3.55$\pm$1.07 11.2% 19.1%
@@
en Large V2 Distil Large V2 4.37$\pm$0.61 3.92$\pm$0.51 3.62$\pm$1.07 10.4% 17.2%
CNN/DailyMail Gemma 7B Gemma 2B 6.54$\pm$0.47 5.84$\pm$0.39 4.37$\pm$0.54 10.6% 33.1%
Qwen 7B Qwen 0.5B 12.01$\pm$1.08 11.20$\pm$1.03 3.34$\pm$0.45 6.8% 72.2%
Llama2 7B Sheared Llama 1.3B 11.33$\pm$0.94 10.69$\pm$0.83 3.65$\pm$0.62 5.7% 67.8%
Llama2 13B Sheared Llama 1.3B 3.99$\pm$0.52 3.59$\pm$1.70 3.26$\pm$0.26 10.1% 18.3%
Xsum Gemma 7B Gemma 2B 6.39$\pm$0.50 5.76$\pm$0.42 4.61$\pm$0.55 9.9% 27.9%
Qwen 7B Qwen 0.5B 11.55$\pm$1.39 10.65$\pm$1.51 3.20$\pm$0.42 7.8% 72.3%
Llama2 7B Sheared Llama 1.3B 4.66$\pm$0.42 4.14$\pm$2.88 3.64$\pm$0.56 11.1% 21.8%
Llama2 13B Sheared Llama 1.3B 4.67$\pm$0.41 4.17$\pm$1.70 3.70$\pm$0.44 10.5% 20.7%
Table 6: Average time and standard deviation spent within the speculative sampling algorithm per decoding step. The column “$\Delta$Prof. Time” measures the relative reduction in average time per decoding step (“Baseline” vs. “Exact” and“Sigmoid”). Scaling constants for sigmoid approximation: $\alpha=-10^{3}$ and $\beta=10^{3}$ for ASR, $\alpha=-10^{4}$ and $\beta=10^{4}$ for summarization.
@@ A.6 Effect of scaling logits
Table 7 shows the impact of various logit scaling factors on performance and profiling time of sigmoid approximation on CV16 and Xsum. The values are computed on a random sample of 10% of each dataset.
Draft/Target Scale ($\alpha,\beta$) WER ()($\downarrow$)( ↓ ) Prof. Time $\Delta$%
@@
Draft/Target Scale ($\alpha,\beta$) ROUGE-1 ()($\uparrow$)( ↑ ) Prof. Time $\Delta$%
@@
A.7 Relation to other optimization methods
Our method is orthogonal to other optimizations of speculative decoding. Whenever speculative sampling is used, our kernel can serve as a drop-in replacement for the standard implementation. For example, our proposed method can be integrated with the recently proposed self-speculative decoding approach Zhang et al. (2024b). Instead of using a separate draft model, self-speculative decoding samples draft tokens by skipping some layers of the target model. Afterwards, it follows the same draft verification and resampling procedure as the original speculative decoding, which can be further accelerated with our optimization method.
Our method is also orthogonal to other approaches for accelerating decoding. For instance, FlashAttention Dao et al. (2022); Dao (2024), which focuses on optimizing the attention computation, can be easily combined with our method to further improve efficiency.
A.8 Overhead caused by resampling
To assess the overhead caused by resampling, we followed Chen et al. (2023a) and computed average acceptance rates of draft tokens for various models on 10% of Xsum. Table 8 includes the acceptance rates using a varying number of draft tokens ($\gamma\in\{3,5,10,15\}$), as well as the average execution time per decoding step of the speculative sampling algorithm. Since our exact optimization method aims to generate the same tokens as the baseline implementation, we expect the acceptance rates to be the same. Table 8 shows that this is indeed the case for all choices of $\gamma$ and model combinations. Table 8 also shows that the acceptance rates with the sigmoid optimization method are often higher than the acceptance rates of the baseline and the exact method. However, these higher acceptance rates, do not have a significant effect on the average execution time. In particular, the acceptance rates of the three methods (sigmoid, exact, and baseline) are similar for the Qwen model combination ($\sim$48with $\gamma=10$), but the sigmoid approximation achieves better execution times than the exact optimization and the baseline. The average execution times provided Table 8 are consistent with the ones presented in 3(a).
Target Model Draft Model Method Acceptance Rate Average Execution Time
$\gamma=3$ $\gamma=5$ $\gamma=10$ $\gamma=15$ $\gamma=3$ $\gamma=5$ $\gamma=10$ $\gamma=15$
Gemma 7B Gemma 2B Sigmoid 56.2% 57.0% 57.1% 56.1% 2.95 ms 3.10 ms 3.09 ms 3.12 ms
Exact 48.5% 48.4% 48.6% 46.4% 4.35 ms 4.36 ms 4.40 ms 4.40 ms
Baseline 48.5% 48.4% 48.6% 46.4% 4.86 ms 4.88 ms 4.90 ms 4.96 ms
Qwen 7B Qwen 0.5B Sigmoid 46.2% 47.8% 48.1% 47.8% 2.92 ms 2.89 ms 2.84 ms 2.83 ms
Exact 46.8% 45.7% 48.4% 48.0% 3.91 ms 3.97 ms 3.88 ms 3.95 ms
Baseline 46.8% 45.7% 48.4% 48.0% 4.45 ms 4.49 ms 4.41 ms 4.48 ms
Llama2 7B Sheared Llama 1.3B Sigmoid 58.4% 59.3% 58.1% 59.4% 3.13 ms 3.26 ms 3.20 ms 3.52 ms
Exact 51.5% 53.7% 56.2% 54.4% 3.41 ms 3.42 ms 3.49 ms 3.78 ms
Baseline 51.5% 53.7% 56.2% 54.4% 3.96 ms 3.95 ms 3.96 ms 4.31 ms
Table 8: Comparison of acceptance rates by optimization type for different models on a random sample of 10% of the Xsum dataset with varying $\gamma$.